A perfectly competitive firm has the cost function TC = 1000 + 2Q + 0.1 Q2. What is the lowest price at which this firm can break even


1
Expert's answer
2021-05-26T13:44:52-0400

First derive MC:


"TC = 1000 + 2Q + 0.1Q^2"


"MC = \frac{\partial TC} {\partial Q} = 2 + 0.2Q"


Then derive ATC:


"ATC = \frac{1000 + 2Q +0.1Q^{2} }{Q}"


"ATC = \frac{1000 }{Q} +2Q +0.1Q^{2}"


Now set: MC = ATC


2 + 0.2Q="\frac{1000 }{Q} +2Q +0.1Q^{2}"


"0.2Q – 0.1Q = \frac{1000 }{Q}"


"0.1Q = \frac{1000 }{Q}"


"0.1Q^2 = 1000"


"Q2 = \frac{1000 }{0.1}"


"Q2 = 10,000"


Take the square root of both sides and find:


Q = 100


We know that the firm produces were Price = MR = MC, so we will derive the lowest price from the MC function:


MC = 2 + 0.2Q


We know Q = 100: Substitute in the equation


MC = 2 + 0.2 (100) = 2 + 20 = 22


Price = 22


Therefore, the lowest price at which the firm can breakeven is 22.